OpenStudy (amyna):

express as a single log: 2log a x - 3log a y =?

OpenStudy (agreene):

remember the rule: \[\log_b\left(\frac{m}{n} \right)= \log_b m - \log_bn\]

OpenStudy (agreene):

Also: \[\log_b m^a = a\log_b m\]